OpenStudy (anonymous):

find the concentration of hydroxide,OH-, ions in a 0.100M solution of HNO2,nitrous acid

OpenStudy (anonymous):

HNO2 -> H+ + NO2- n(H+) = n(HNO2) so c(H+) = c(HNO2) pH = - log [H+] pH = 1 pOH = 14-pH pOH = 13 pOH = -log [OH-] 10^-pOH = [OH-] 10^-13 = [OH-] [OH-] = 1*10^-13 M